OpenStudy (anonymous):

A rectangle has a base of 5 inches and a height of 7 inches. If the dimensions are cut in half, what will happen to the area of the rectangle? possible answers: It will be half of the original size. It will be the same as the original size. It will be four times the original size. It will be one-fourth the original size.

OpenStudy (anonymous):

It will be one fourth

OpenStudy (lgbasallote):

area is base x height if the dimensions would be cut then it would be \[\large \frac{5}{2} \times \frac{7}{2}\] what does that equal?